# OWLed: Outlier-weighed Layerwise Pruning for Efficient Autonomous Driving Framework

### v2

@@ Abstract
The integration of Large Language Models (LLMs) into autonomous driving systems offers promising enhancements in environmental understanding and decision-making. However, the substantial computational demands of deploying LLMs locally on vehicles render this approach unfeasible for real-world automotive applications. To address this challenge, we introduce OWLed, the Outlier-weighed Layerwise Pruning for Efficient Autonomous Driving Framework that leverages outlier-weighted layerwise sparsity for model compression. Our method assigns non-uniform sparsity ratios to different layers based on the distribution of outlier features, significantly reducing the model size without the need for fine-tuning. To ensure the compressed model adapts well to autonomous driving tasks, we incorporate driving environment data into both the calibration and pruning processes. Our empirical studies reveal that the encoder component is more sensitive to pruning than the LLM, highlighting its critical role in the system. Experimental results demonstrate that OWLed outperforms existing methods in perception, action prediction, and language understanding while substantially lowering computational requirements. These findings underscore the potential of combining advanced pruning techniques with LLMs to develop efficient and robust autonomous driving systems capable of handling complex scenarios. Code is available at https://github.com/JiaxiLi1/OWLed.
@@ I Introduction
Autonomous driving technology has made significant progress in recent years, with an increasing number of autonomous vehicles being tested and deployed on public roads [1, 2]. There are generally two types of autonomous driving systems: one is a modular design with different sub-modules to complete various tasks such as perception, prediction, and planning [3, 4, 5, 6]. Such a decoupled design may lead to cumulative errors as the module cannot access the sensor data. The other is an end-to-end design that directly inputs sensor data into a series of neural networks to obtain control signals [7, 8, 9]. However, systems built with traditional neural networks often struggle to handle long-tail scenarios and complex urban environments, limiting their ability to process the diverse and unpredictable situations encountered in real-world driving [1, 2, 10].
Recently, large language models (LLMs) have exhibited capabilities approaching Artificial General Intelligence, including common sense understanding, reasoning, and knowledge retrieval [11, 12, 13]. LLMs’ remarkable capabilities make them a promising solution for addressing the aforementioned challenges in autonomous driving systems. This has sparked several attempts to integrate LLMs into autonomous vehicles to enhance understanding of the driving environment, explain perception results [14], generate textual driving decisions [15], and translate those decisions into executable commands [10]. These works primarily fall into two categories: The first category leverages pre-trained LLMs, such as ChatGPT, through API interfaces to process sensor data (i.e., camera and radar inputs) for environmental understanding and decision-making [14, 16, 17, 15, 18]. This approach can fully utilize state-of-the-art pre-trained LLMs to achieve optimal performance. However, its drawback lies in its dependence on network conditions and the data processing speed of servers, which limits its effectiveness in complex driving scenarios [10]. The second category involves training and deploying LLMs locally on autonomous vehicle systems. This method could address the autonomous driving system’s requirements for cloud computing resources and network communication quality [19, 20, 9, 21, 22, 23, 24]. However, as LLMs typically operate at the billion-parameter scale, this approach places substantial demands on local computational resources, thereby limiting their practical applications in autonomous vehicles.
@@
Recent advancements have explored pruning LLMs without fine-tuning. These studies have demonstrated that a substantial portion of the model’s parameters can be eliminated in a single step while incurring only minimal performance degradation. For example, sparseGPT [36] is a one-shot pruning method based on layer-wise reconstruction. It uses the Optimal Brain Surgeon (OBS) algorithm to determine which weights to prune. Wanda [37], on the other hand, is a simple yet effective pruning approach that considers both weight magnitudes and input activations. Wanda is computationally more efficient than SparseGPT as it doesn’t require the approximation of second-order information. Note that both SparseGPT and Wanda are designed to apply uniform layerwise sparsity when pruning LLMs. This means that each layer in the model is pruned to the same degree of sparsity, regardless of its position or function within the network. More recently, research has shown that non-uniform layerwise sparsity can potentially yield better performance [38]. Building on this insight, Yin et al. [38] proposed OWL as a non-uniform layerwise sparsity strategy based on the distribution of outlier features. OWL captures the impact of outlier features in LLMs and assigns sparsity ratios proportional to each layer’s outlier ratio. OWL has shown demonstrated superior performance compared to uniform pruning methods, particularly at high sparsity levels. Importantly, it can be combined with existing pruning techniques to further improve performance.
While many attempts have been made to leverage LLMs in autonomous driving, the research of computationally efficient LLMs for end-to-end autonomous driving is still at the very preliminary stage. To address the computational challenges associated with local LLM deployment in autonomous vehicles, we propose the first effective LLM-assisted autonomous driving framework by utilizing outlier-weighted layerwise sparsity to compress LLMs, termed OWLed. To ensure effective adaptation of LLMs to the specific demands of autonomous driving, we incorporate downstream task data into both the calibration process and the optimization of the pruning procedure. This tailored approach allows for more precise and context-specific model compression. The main contributions of this paper are summarized as follows:
  1. We propose an effective LLM assisted autonomous driving framework that significantly enhances computational efficiency while maintaining the powerful reasoning capabilities of LLMs. By employing outlier-weighted layerwise sparsity, our approach effectively addresses the computational constraints inherent in deploying LLMs in vehicles. This framework is designed to process and reason over complex environmental data from multiple sensory modalities and natural language inputs, enabling more sophisticated and efficient decision-making in autonomous vehicles.
@@ II-B Neural network pruning
Traditional LLM pruning methods usually require fine-tuning or retraining to restore performance. For example, LLM-Pruner is proposed in [42] for structural pruning. It uses Taylor pruning to remove entire weight rows based on their impact on the model output. Recent research has pivoted towards developing LLM pruning methods that do not require fine-tuning. SparseGPT [36] uses the Optimal Brain Surgeon (OBS) algorithm to select weights for pruning and then update the remaining weights to minimize output reconstruction error. In this way, it addresses the challenge of computing the Hessian inverse. Wanda [37] calculates an importance score for each weight as the product of the weight magnitude and the $\ell$2 norm of all input features connected to that weight. It then ranks and prunes weights based on this importance score, without relying on second-order information or weight updates. Dynamic sparsity [43, 44] is extended in [45] to efficiently fine-tune sparse LLM without weight updating. The JunK DNA hypothesis [46] investigates optimal overall sparsity for different downstream tasks. However, most existing methods apply uniform sparsity across all layers, overlooking the presence of outlier features in LLMs. Inspired by this, OWL [38] introduces a non-uniform layerwise sparsity strategy. This approach calculates each layer’s Layerwise Outlier Distribution (LOD) and assigns sparsity ratios proportional to these outliers, demonstrating promising performance gain.
Inspired by OWL, we leverage outlier information to compute optimal sparsity within an autonomous driving framework. To the best of our knowledge, this is the first work to integrate LLMs with autonomous driving systems in the context of pruning for efficiency.
@@ III OWLed: Outlier-weighed Layerwise Pruning for Efficient Autonomous Driving Framework
Figure 1: Illustration of the architecture of OWLed.
1:  Input: $A$: Calibration data, $M$: Threshold coefficient, $\lambda$: Limiting value, $L$: Number of layers;
@@
6:  Calculate target sparsity $[S_{1},S_{2},...,S_{L}]$ by using $S_{i}\propto 1-D_{i}$;
@@ III-B Layerwise Outlier Distribution-based pruning
We first calculate the LOD of each layer of the model. For a model with $L$ layers, we calculate LOD = $[D_{1},D_{2},...,D_{L}]$. Specifically, for the $\ell$th layer, we calculate its outlier ratio $D_{\ell}$ as follows:
@@
where $A^{\ell}_{ij}=\|X_{j}\|_{2}\cdot|W^{\ell}_{ij}|$ is the outlier score of the weight $W^{\ell}_{ij}$, calculated as the $\ell_{2}$ norm of all input features connected to this weight multiplied by the absolute value of the weight. $\bar{A^{\ell}}$ is the average of all outlier scores of the layer, $M$ is a hyperparameter used to determine the threshold of the outlier, and $I(\cdot)$ denotes the indicator function, which returns 1 when the condition is met and 0 otherwise.
@@
Given the target model sparsity $S$, we determine the sparsity $[S_{1},S_{2},...,S_{L}]$ of each layer, based on the principle that layers with higher outlier ratios should have lower sparsity. Specifically, we set $S_{i}\propto 1-D_{i}$. To prevent the sparsity difference between layers from being too large, we introduce a hyperparameter $\lambda$ to limit the sparsity of each layer to the range of $[S-\lambda,S+\lambda]$, while keeping the average sparsity of all layers as $S$.
@@
In[38], the LOD is calculated using general text data from the C4 dataset [47] as calibration data. However, in our autonomous driving context, the input data consists of vector representations of driving scenarios, which significantly differs from general text data. This disparity in input data can lead to substantially different LOD calculations, potentially affecting the pruning results. To better align with our downstream task requirements and to leverage task-specific information, we propose to use our driving scenario vector data to calculate the LOD values for pruning. This approach ensures that the pruning process is tailored to the specific needs and characteristics of our autonomous driving task, potentially leading to more effective pruning results.
Figure 2: The evaluation results of perception and action prediction achieved by different pruning techniques.
Sparsity_ratio Method GPT Grading $\uparrow$
0.3 Magnitud 7.03
@@
TABLE I: GPT Grading of the driving QA outputs.
IV Experiments
IV-A Experiment setup
In this section, we verify the performance of OWLed. We employ the driving QA dataset in [9] as our evaluation dataset. The dataset comprises three subsets: a 100k dataset for vector encoder pretraining, a 10k dataset for fine-tuning, and a 1k dataset for evaluation. The data is collected in a 2D simulated driving environment with randomized road conditions.
We adopt the evaluation metrics in [9] to evaluate the proposed model’s perception and action prediction ability. These metrics include mean absolute error for the number of cars prediction ($E_{car}$) and pedestrians prediction ($E_{ped}$), accuracy of traffic light detection ($ACC_{TL}$), mean absolute distance error in meters for traffic light distance prediction ($D_{TL}$) and normalized steering wheel angle ($E_{lat}$), and weighted cross-entropy loss for token prediction on the evaluation set ($L_{token}$). Additionally, we utilize GPT-3.5 to evaluate the quality of the model’s responses about the driving environments. This technique allows for rapid and consistent assessment of answers without fixed solutions [48, 49, 50]. Specifically, we provide GPT-3.5 with the language-based observation description used during dataset generation, test questions, and the model’s answer. Based on the given observations, the GPT then generates an assessment and a score from 0 to 10 for each response. The results of the GPT evaluations are the model’s final score averaged across all questions.
Sparsity_ratio Method $E_{car}\downarrow$ $E_{ped}\downarrow$ $L_{token}\downarrow$ $ACC_{TL}\uparrow$ $D_{TL}\downarrow$ $E_{lat}\downarrow$
@@ IV-A Experiment setup
TABLE II: Evaluation of pruning encoder ablation.
@@ IV-B1 Baselines
[…]
@@ IV-B2 Results
Wanda and OWL demonstrate similar performance. Both methods consider input activation when calculating weight outlier values for pruning. Additionally, OWL employs non-uniform sparsity pruning by calculating the LOD. However, OWL’s performance could be influenced by its hyperparameter, which leads to its effectiveness can not be guaranteed across different tasks. For example, while OWL achieves comparable performance in terms of $ACC_{TL}$, it fails to maintain its efficiency for the steering wheel angle prediction, as indicated by $E_{lat}$.

### v3

@@ Abstract
The integration of Large Language Models (LLMs) into autonomous driving systems offers promising enhancements in environmental understanding and decision-making. However, the substantial computational demands of deploying LLMs locally on vehicles render this approach unfeasible for real-world automotive applications. To address this challenge, we introduce OWLed, the Outlier-weighed Layerwise Pruning for Efficient Autonomous Driving Framework that leverages outlier-weighed layerwise sparsity for model compression. Our method assigns non-uniform sparsity ratios to different layers based on the distribution of outlier features, significantly reducing the model size without the need for fine-tuning. To ensure the compressed model adapts well to autonomous driving tasks, we incorporate driving environment data into both the calibration and pruning processes. Our empirical studies reveal that the encoder component is more sensitive to pruning than the LLM, highlighting its critical role in the system. Experimental results demonstrate that OWLed outperforms existing methods in perception, action prediction, and language understanding while substantially lowering computational requirements. These findings underscore the potential of combining advanced pruning techniques with LLMs to develop efficient and robust autonomous driving systems capable of handling complex scenarios. Code is available at https://github.com/JiaxiLi1/OWLed.
@@ I Introduction
Autonomous driving technology has made significant progress in recent years, with an increasing number of autonomous vehicles being tested and deployed on public roads [1, 2]. There are generally two types of autonomous driving systems: one is a modular design with different sub-modules to complete various tasks such as perception, prediction, and planning [3, 4, 5, 6]. While this modular approach provides system interpretability, it suffers from error accumulation and information loss between components because its modules cannot directly access sensor data. The other is an end-to-end design that directly inputs sensor data into a series of neural networks to obtain control signals [7, 8, 9]. However, systems built with traditional neural networks often struggle to handle long-tail scenarios and complex urban environments, limiting their ability to process the diverse and unpredictable situations encountered in real-world driving [1, 2, 10].
Recently, large language models (LLMs) have exhibited capabilities approaching Artificial General Intelligence, including common sense understanding, reasoning, and knowledge retrieval [11, 12, 13]. LLMs’ remarkable capabilities make them a promising solution for addressing the aforementioned challenges in autonomous driving systems. This has sparked several attempts to integrate LLMs into autonomous vehicles to enhance understanding of the driving environment, explain perception results [14], generate textual driving decisions [15], and translate those decisions into executable commands [10]. These works primarily fall into two categories: The first category leverages pre-trained LLMs, such as ChatGPT, through API interfaces to process sensor data (i.e., camera and radar inputs) for environmental understanding and decision-making [14, 16, 17, 15, 18]. However, its drawback lies in its dependence on network conditions and the data processing speed of servers, which limits its effectiveness in complex driving scenarios [10]. The second category involves training and deploying LLMs locally on autonomous vehicle systems. This method could address the autonomous driving system’s requirements for cloud computing resources and network communication quality [19, 20, 9, 21, 22, 23, 24]. However, as LLMs typically operate at the billion-parameter scale, this approach places substantial demands on local computational resources, thereby limiting their practical applications in autonomous vehicles.
@@
Recent advancements have explored pruning LLMs without fine-tuning. These studies have demonstrated that a substantial portion of the model’s parameters can be eliminated in a single step while incurring only minimal performance degradation. For example, sparseGPT [36] is a one-shot pruning method based on layer-wise reconstruction. It uses the Optimal Brain Surgeon (OBS) algorithm to determine which weights to be pruned. Wanda [37], on the other hand, is a simple yet effective pruning approach that considers both weight magnitudes and input activations. Wanda is computationally more efficient than SparseGPT as it doesn’t require the approximation of second-order information. Note that both SparseGPT and Wanda are designed to apply uniform layerwise sparsity when pruning LLMs. This means that each layer in the model is pruned to the same degree of sparsity, regardless of its position or function within the network. More recently, research has shown that non-uniform layerwise sparsity can potentially yield better performance [38]. Building on this insight, Yin et al. [38] proposed OWL as a non-uniform layerwise sparsity strategy based on the distribution of outlier features. OWL captures the impact of outlier features in LLMs and assigns sparsity ratios proportional to each layer’s outlier ratio. OWL has shown demonstrated superior performance compared to uniform pruning methods, particularly at high sparsity levels. Importantly, it can be combined with existing pruning techniques to further improve performance.
While many attempts have been made to leverage LLMs in autonomous driving, the research of computationally efficient LLMs for end-to-end autonomous driving is still at the very preliminary stage. To address the computational challenges associated with local LLM deployment in autonomous vehicles, we propose the first effective LLM-assisted autonomous driving framework by utilizing outlier-weighed layerwise sparsity to compress LLMs, termed OWLed. To ensure effective adaptation of LLMs to the specific demands of autonomous driving, we incorporate downstream task data into both the calibration process and the optimization of the pruning procedure. This tailored approach allows for more precise and context-specific model compression. The main contributions of this paper are summarized as follows:
  1. We propose an effective LLM assisted autonomous driving framework that significantly enhances computational efficiency while maintaining the powerful reasoning capabilities of LLMs. By employing outlier-weighed layerwise sparsity, our approach effectively addresses the computational constraints inherent in deploying LLMs in vehicles. This framework is designed to process and reason over complex environmental data from multiple sensory modalities and natural language inputs, enabling more sophisticated and efficient decision-making in autonomous vehicles.
@@ II-B Neural network pruning
Traditional LLM pruning methods usually require fine-tuning or retraining to maintain performance. For example, LLM-Pruner is proposed in [42] for structural pruning. It uses Taylor pruning to remove entire weight rows based on their impact on the model output. Recent research has pivoted towards developing LLM pruning methods that do not require fine-tuning. SparseGPT [36] uses the OBS algorithm to select weights for pruning and then updates the remaining weights to minimize output reconstruction error. In this way, it addresses the challenge of computing the Hessian inverse. Wanda [37] evaluates an importance score for each weight by calculating the product of weight magnitude and the $\ell$2 norm of associated input features. Accordingly, weights can be pruned based on the importance score, without relying on second-order information or weight updates. Dynamic sparsity [43, 44] is extended in [45] to efficiently fine-tune sparse LLM without weight updating, while the JunK DNA hypothesis [46] investigates optimal overall sparsity for different downstream tasks. However, most existing methods apply uniform sparsity across all layers, overlooking the presence of outlier features in LLMs. Inspired by this, OWL [38] introduces a non-uniform layerwise sparsity strategy. This approach calculates each layer’s Layerwise Outlier Distribution (LOD) and assigns sparsity ratios proportional to these outliers, demonstrating promising performance gain.
Inspired by OWL, we leverage outlier information to compute optimal sparsity within an autonomous driving framework. Our method utilizes the insight that different layers contribute differently to the model’s decision-making process in autonomous driving scenarios. By identifying and preserving crucial outlier features specific to driving tasks, OWLed maintains high performance even at high sparsity levels. To the best of our knowledge, this is the first work to integrate LLMs with autonomous driving systems in the context of pruning for efficiency.
@@ III OWLed: Outlier-weighed Layerwise Pruning for Efficient Autonomous Driving Framework
Figure 1: Illustration of the architecture of OWLed and the LOD-based pruning. The autonomous driving model consists of a vector encoder for processing multimodal input data and an LLM for decision making. The model uses outlier-weighed layer-wise pruning and incorporates input activation to achieve efficient compression while maintaining the model performance.
1:Input: AA: Calibration data, $M$: Threshold coefficient, $\lambda$: Limiting value, $L$: Number of layers;
@@
6: Calculate target sparsity $S_{1},S_{2},...,S_{L}]$ by using $S_{i}\propto 1-D_{i}$;
@@ III-B Layerwise Outlier Distribution-based pruning
[…]
@@
where $A^{\ell}_{ij}=\|X_{j}\|_{2}\cdot|W^{\ell}_{ij}|$ is the outlier score of the weight $W^{\ell}_{ij}$, calculated as the $\ell_{2}$ norm of all input features connected to this weight multiplied by the absolute value of the weight. $\bar{A^{\ell}}$ is the average of all outlier scores of the layer, $M$ is a hyperparameter used to determine the threshold of the outlier, which is determined through grid search on the validation set (see Section IV for details). $I(\cdot)$ denotes the indicator function, which returns 1 when the condition is met and 0 otherwise. $C_{in}$ and $C_{out}$ denote the input and output channel dimensions of the layer, respectively.
@@
Given the target model sparsity $S$, we determine the sparsity $S_{1},S_{2},...,S_{L}]$ of each layer, based on the principle that layers with higher outlier ratios should have lower sparsity. Specifically, we set $S_{i}\propto 1-D_{i}$. To prevent the sparsity difference between layers from being too large, we introduce a hyperparameter $\lambda$ to limit the sparsity of each layer to the range of $S-\lambda,S+\lambda]$, while keeping the average sparsity of all layers as $S$.
III-C Calibration data for OWLed
Traditional pruning methods [38, 37, 36] utilize random text snippets from the C4 datasets, which contains general linguistic patterns but lacks driving-specific semantics.
@@ III-C Calibration data for OWLed
Regarding the calculation of LOD, in[38], the LOD is calculated using data from the C4 dataset [47] as calibration data. However, in our autonomous driving context, the input data consists of vector representations of driving scenarios, which significantly differs from general text data. This disparity in input data can lead to substantially different LOD calculations, potentially affecting the pruning results. To better align with our downstream task requirements and to leverage task-specific information, we propose to use our driving scenario vector data to calculate the LOD values for pruning. This approach ensures that the pruning process is tailored to the specific needs and characteristics of our autonomous driving task, potentially leading to more effective pruning results.
Figure 2: The evaluation results of perception and action prediction achieved by different pruning techniques. Results show performance across multiple metrics including car detection error ($E_{car}$), pedestrian detection error ($E_{ped}$), token prediction loss ($L_{token}$), traffic light detection accuracy ($ACC_{TL}$), traffic light distance error ($D_{TL}$), and lateral control error ($E_{lat}$) at sparsity ratios varying from 0.1 to 0.5.
IV Experiments
IV-A Experiment setup
In this section, we verify the performance of OWLed. We explore two types of calibration data for calculating LOD. Traditional pruning methods typically use general text data from the C4 dataset [47], which consists of a diverse range of web-crawled text. In contrast, we propose using domain-specific driving scenario data collected from a custom-built 2D simulator [9]. This dataset includes continuous driving data from 15 diverse virtual environments with randomly generated traffic conditions, containing vector representations of driving scenarios that are directly relevant to our autonomous driving tasks. The use of task-specific calibration data allows for more targeted and effective pruning in the context of autonomous driving.
We employ the driving QA dataset in [9] as our evaluation dataset. The dataset comprises three subsets: a 100k dataset for vector encoder pretraining, a 10k dataset for fine-tuning, and a 1k dataset for evaluation.
We adopt the evaluation metrics in [9] to evaluate the proposed model’s perception and action prediction ability. These metrics include mean absolute error for the number of cars prediction ($E_{car}$) and pedestrians prediction ($E_{ped}$), accuracy of traffic light detection ($ACC_{TL}$), mean absolute distance error in meters for traffic light distance prediction ($D_{TL}$) and normalized steering wheel angle ($E_{lat}$), and weighted cross-entropy loss for token prediction on the evaluation set ($L_{token}$). Additionally, we utilize GPT-3.5 to evaluate the quality of the model’s responses to the driving environments. This technique allows for rapid and consistent assessment of answers without fixed solutions [48, 49, 50]. Specifically, we provide GPT-3.5 with the language-based observation description used during dataset generation, test questions, and the model’s answer. Based on the given observations, the GPT then generates an assessment and a score from 0 to 10 for each response. The results of the GPT evaluations are the model’s final score averaged across all questions.
Sparsity ratio Method GPT Grading $\uparrow$
0.3 Magnitude 7.03
@@ IV-A Experiment setup
TABLE I: GPT grading results of the driving QA outputs achieved by different pruning methods at sparsity ratios of 0.3 and 0.4, and baseline performances achieved by the dense model, constant answers, random answers, and GPT-generated answers. Higher scores indicate better quality of responses to driving-related questions.
Sparsity ratio Method $E_{car}\downarrow$ $E_{ped}\downarrow$ $L_{token}\downarrow$ $ACC_{TL}\uparrow$ $D_{TL}\downarrow$ $E_{lat}\downarrow$
@@
TABLE II: Evaluation of pruning encoder ablation. Comparison between OWLed variants (OWLed, OWLed-separate, OWLed-global) across different metrics at sparsity ratios of 0.3 and 0.4, respectively.
@@ IV-B1 Baselines
The original autonomous driving model from [9] is used as one of our baselines, referred to as Dense. Additionally, we create three more baselines by adopting different state-of-the-art pruning techniques to this architecture, i.e., Magnitude [51], Wanda [37] and OWL [38]. We test all the methods under comparison with different sparsity levels, i.e., $sparsityratio\in(10\%,20\%,30\%,40\%,50\%)$, and set the number of calibration data as 128 for all baselines. For the OWL implementation, we conduct a comprehensive grid search to optimize the hyperparameters, selecting $\lambda=0.1$ and $M=5$ from candidate ranges of $\lambda\in(\text{0.02,0.05,0.08,0.1,0.2})$ and $M\in(\text{3, 5, 7, 10})$, following [38].
@@ IV-B2 Results
Wanda and OWL demonstrate similar performance. Both methods consider input activation when calculating weight outlier values for pruning. Additionally, OWL employs non-uniform sparsity pruning by calculating the LOD. However, OWL’s performance could be influenced by its hyperparameter, which leads to its effectiveness can not be guaranteed across different tasks. For example, while OWL achieves comparable performance in terms of $ACC_{TL}$, it fails to maintain its efficiency for the steering wheel angle prediction, as indicated by $E_{lat}$. We note that for the $ACC_{TL}$ metric, OWLed even outperforms the dense model at 0.2 sparsity, which suggests that the dense model is not always better than the sparse model. This could be because of the sparse model’s ability to prevent overfitting by decreasing the capacity of the model [52, 53].